OpenStudy (anonymous):

need to calculate the upper Reimann sum for g(x) x^4,,,,anyone?? Not sure I am even asking the question correctly....I have 1. Consider g(x) = x^4 on the interval [0,1]...a) calculate the Upper Reimann Sum U=(f,Psub n) where Psub n = {j/n}

OpenStudy (anonymous):

What is the interval in question? :-)

OpenStudy (anonymous):

...where Psub n = {j/nI j= 0,1,2,33..., n} and we were given the formula for the sum....

OpenStudy (anonymous):

....sigma j=1 to n j^4 = n9n+1)(2n+1)(3n^2 + 3n -1) /30

OpenStudy (anonymous):

Not sure where I am supposed to start..

OpenStudy (anonymous):

bmp....are you there?

OpenStudy (turingtest):

I'm having a hard time understanding your question you want to approximate the integral with of g(x)=x^4 in the interval [0,1] with a reimann sum, correct?

OpenStudy (anonymous):

Sorry, got distracted. But @TuringTest can pick up from here :-)

OpenStudy (anonymous):

thanks!

OpenStudy (turingtest):

how many pieces do you want to break it up into are you told to use a particular value for n?

OpenStudy (anonymous):

yes, think he wants us to do the long way with the Upper Reimann Sum...

OpenStudy (anonymous):

ahm...I eidited my question, but don't see it all there.......here goes again...

OpenStudy (anonymous):

Psub n ={j/nI j=0, 1, 2, ..., n} does this tell be how many pieces to use, etc?

OpenStudy (anonymous):

would this just result in the summation formula he gave ? n(n+1)(2n + 1)(3n^2 + 3n -1) /30?

OpenStudy (turingtest):

\[\int_0^1x^4dx\approx\sum_{j=1}^nf(\Delta x^*)\Delta x\]it doewsn't say how many pieces to use, so I suppose they want n to go to infinity, which is the definition of the integral and yeah I think it will just wind up being that formula, but you need to make sure to keep \(\Delta x\) in there

OpenStudy (anonymous):

yes, I see that he says j=1 so height is a function of x^4 and each rect. has width 1/n

OpenStudy (anonymous):

ok, I did this same thing then a couple of weeks ago for x^3...but can you continue to help me make sure I'm correct?? :-)

OpenStudy (turingtest):

\[\int_0^1x^4dx=\lim_{n\to\infty}\sum_{j=1}^nf(\Delta x_j^*)\Delta x=\lim_{n\to\infty}\sum_{j=1}^n\frac{j^4}{n^4}\cdot\frac1n=\lim_{n\to\infty}\sum_{j=1}^n\frac{j^4}{n^6}\]and now we get to use that hideous formula of yours for the sum of j^4

OpenStudy (turingtest):

typo*\[\int_0^1x^4dx=\lim_{n\to\infty}\sum_{j=1}^nf(\Delta x_j^*)\Delta x=\lim_{n\to\infty}\sum_{j=1}^n\frac{j^4}{n^4}\cdot\frac1n=\lim_{n\to\infty}\sum_{j=1}^n\frac{j^4}{n^5}\]

OpenStudy (anonymous):

last time I subbed in i/n^3 because ?? my summation ended up like 1/4 summation j=1 to n j^3??

OpenStudy (turingtest):

\[\lim_{n\to\infty}\frac1{n^5}\sum_{j=1}^ni^4=\lim_{n\to\infty}\frac1{n^5}{n(n+1)(2n+1)(3n^2 + 3n -1)\over30}\]simplify all that nastiness on the right

OpenStudy (turingtest):

where did you get cubed from? I thought we were doing ^4th power

OpenStudy (anonymous):

ok, it's a 1/n^5 outside the summation and NOT just the fraction 1/5? (the cubed is from the problem I did a few weeks ago that muxt have been simial...

OpenStudy (turingtest):

yeah we can take the whole n outside the summation sign because it is constant with respect to the sum it still has to stay inside the limit sign though

OpenStudy (turingtest):

in this case the n term is 1/n^5 now we only need to sum j, for which we can use your formula

OpenStudy (anonymous):

ok, in the earlier problem I had a 1/4 out front of the summation but started simpl,ifying with a 1/n^4

OpenStudy (turingtest):

well a lot of stuff should cancel now comes the awful algebra part where you need to expand the numerator

OpenStudy (anonymous):

ok, I must have just left the n in n^4 out and the prof didn't catch it!

OpenStudy (anonymous):

yes, it did in the last formula he gave us for n^3...simpl;ied down with the 1/n^4 that was out front

OpenStudy (zarkon):

just loot at the highest power of n in the numerator...\[n\cdot n\cdot 2n\cdot3n^2\]

OpenStudy (zarkon):

*look

OpenStudy (anonymous):

let me see what the prof gave us again...maybe I mistyped...

OpenStudy (anonymous):

n(n+1)(2n+1)(3n^2+3n_1) all over 30

OpenStudy (turingtest):

@Zarkon oh, so we don't have to expand after all :) as usual: Hail Zarkon!

OpenStudy (zarkon):

never work harder than you have too ;)

OpenStudy (turingtest):

of course. work smarter, not harder

OpenStudy (anonymous):

yes...we can cancel some things.....the prof didn't expand the last portion for the previous prob either, I now remember....

OpenStudy (anonymous):

oh, thanks guys!! think I have it...now I just have to evealuate a few values in....THEN use Leibnitz'value twice to integrate HIS way.... lvoely! I may be getting back to you later!!!! :-):-)

OpenStudy (turingtest):

welcome, good luck :)